### v1

Abstract
Trade restrictions, the COVID-19 pandemic, and geopolitical conflicts has significantly exposed vulnerabilities within traditional global supply chains. These events underscore the need for organisations to establish more resilient and flexible supply chains. To address these challenges, the concept of the autonomous supply chain (ASC), characterised by predictive and self-decision-making capabilities, has recently emerged as promising solution. However, research on ASCs is relatively limited, with no existing studies on their implementations. This paper aims to address this gap by presenting an implementation of ASC using a multi-agent approach. It proposes a methodology for the analysis and design of such an agent-based ASC system (A2SC). This paper provides a concrete case study, the autonomous meat supply chain, which showcases the practical implementation of the A2SC system using the proposed methodology. Additionally, a system architecture and a toolkit for developing A2SC systems are presented. Despite with limitations, this paper demonstrates a promising approach for implementing an effective ASC system.
keywords:
Autonomous supply chain, Multi-agent system, Autonomous agents, Perishable foods
[…]
[inst1]organization=Institute for Manufacturing, Department of Engineering, University of Cambridge, addressline=17 Charles Babbage Road, city=Cambridge, postcode=CB3 0FS, country=United Kingdom
[…]
[inst1]organization=Fetch.ai, addressline=St John’s Innovation Centre, Cowley Rd, city=Cambridge, postcode=CB4 0WS, country=United Kingdom
1 Introduction
Trade restrictions, the COVID-19 pandemic, and geopolitical conflicts has seriously exposed vulnerabilities within traditional global supply chains [Handfield et al., 2020, Shih, 2020], highlighting the need for organisations to establish more resilient and flexible supply chains. Indeed, traditional supply chains exhibit limited agility and resilience, with over 70% of surveyed supply chain leaders acknowledging the need for enhanced agility and operational resilience [NelsonHall, 2021]. The model of autonomous supply chains (ASC) has been proposed as a concept solution to achieve supply chain resilience and structural flexibility [Xu et al., 2022].
Traditional supply chains still heavily rely on manual processes, which can often introduce communication delays and human errors, ultimately leading to increased costs [NelsonHall, 2021]. Intelligent digital technologies, such as AI and robotics, are penetrating into supply chain domain, automating various processes, particularly those error-prone and routine tasks. However, these automation efforts mainly focuses on individual processes, with limited integration between processes, let alone inter-firm integration. The needs for better supply chain performance require more horizontal and vertical integration [Frohlich and Westbrook, 2001, Stevens and Johnson, 2016]. ASCs emphases interconnection, integration, automation, and self-decision-making. Within this framework, the supply chain has capacity to predict disruptions and automatically adapt to changes through reconfiguration and adjustments. Consequently, ASCs offer a promising solution to address challenges in an uncertain and turbulent business environment.
The concept of supply chains equipped with such predictive and self-adaptive capabilities have been a topic of discussion for many years. A “smarter” supply chain, capable of autonomous learning and decision-making without human intervention, was put forth by Butner [2010]. This envisioned supply chain was characterised as instrumented, interconnected, and intelligent. Furthermore, Wu et al. [2016] conducted a literature review on the impact of digital technologies on SCM and proposed three additional defining characteristics for this emerging supply chain: automation, integration, and innovation. Expanding on these ideas, Calatayud et al. [2019] proposed the concept of the self-thinking supply chain, which is distinguished by autonomous and predictive capabilities relying on IoT and AI-enabled connectivity. Moreover, Xu et al. [2022] extended these concepts by formally proposing the autonomous supply chain model. This work extensively delved into the theoretical aspects of ASCs, including their formal definition, defining characteristics, and a set of auxiliary concepts. Instead of solely outlining the characteristics of this intelligent supply chain, it presented a five-layer conceptual model and a seven-level supply chain autonomy reference model. These models serve as practical guidelines for the implementation of ASC systems.
However, despite the existence of initial concepts and theoretical studies in literature, the development of ASCs remains in its nascent stages. Notably, none of these works have extensively addressed the technical aspects of ASCs. This paper, therefore, aims to bridge this gap by focusing on the technical implementation of ASCs. While Xu et al. [2022] presented a short high-level overview of an ASC implementation case, this paper delves comprehensively into all facets of implementing an ASC system, employing a Multi-Agent System (MAS) approach.
A supply chain consists of a network of business entities engaged in various activities that transform raw materials into finished products and distribute them to end customers [Christopher, 2016]. These activities include demand forecast, material procurement, inventory management, and order fulfilment. Effectively managing the flow of goods and its associated information and financial flows is crucial in supply chain management (SCM). To achieve this automatically, interoperability between independent and geographically distributed entities must be facilitated. MAS approaches, which achieved distributed problem-solving through multiple intelligent agents interacting within a shared environment, are inherently well-suited for architecturally modelling supply chains [Wooldridge and Jennings, 1995, Fox et al., 2001, Xu et al., 2023]. These agents can be seen as an independent, goal-oriented software process that operates autonomously, communicating and coordinating with other agents as needed. This paper investigates the creation of an agent-based ASC (A2SC) system that facilitates interoperability and automation among distributed heterogeneous supply chain entities.
As described in Xu et al. [2022], entirely automating a supply chain involves replacing manual processes by automated ones and establishing the connections between these automated processes. These automated processes, which are constituent building blocks for ASCs, can be managed through a set of structural entities defined in Xu et al. [2022]. Moreover, these structural entities are responsible for managing the three flows that traverse these interconnected processes. In A2SCs, these structural entities can be represented by software agents, autonomously managing its internal entities and interacting with external entities.
Despite agent-based technology being proposed for use in SCM for over two decades, it current technical readiness level (TRL) remains relatively low, and its industrial adoption is limited [Xu et al., 2021]. This paper aims to investigate fundamental issues relevant to the design and development of a practical A2SC system, including its analysis, design, and implementation.
The main contributions of this paper can be summarised as follows:
  1. (1)
    We investigate the application of the MAS approach in supply chain domain and introduce an Agent-based Autonomous Supply Chain (A2SC).
  2. (2)
    We present a methodology designed for the analysis and design of A2SC systems, and detail the realisation of A2Sc systems using this methodology.
  3. (3)
    We present a detailed concrete case study, the autonomous meat supply chain, demonstrating the practical implementation of an A2SC system.
  4. (4)
    We report a system architecture and a toolkit for A2SC prototype development, potential offering guidance for similar decentralised systems.
The rest of this paper is organised as follows. Section 2 reviews related work. Section 3 presents agent-based autonomous supply chains and a methodology for its analysis and design. Section 4 details an A2SC implementation case study. Section 5 discusses the limitation and implications of this work. Finally, Sections 6 concludes this paper and discusses future work.
2 Related Work
Given that ASCs remain at its early stages, tehre are limited existing literature related to their technical implementation. This section reviews the work related to the use of MAS approaches in SCM, which is relevant for implementing A2SCs.
The adoption of MAS approaches in SCM can be traced back to the development of the Integrated Supply Chain Management System (ISCM) at the University of Toronto in the early 1990s [Fox et al., 1993]. The ISCM was composed of a collection of software agents, with each responsible for one or more activities within the supply chain. These agents coordinated with one another to plan and execute supply chain functions. Fox et al. [2001] investigated key issues and provided detailed solutions for constructing such an agent-based supply chain architecture. Similarly, Swaminathan et al. [1998], another early representative work, employed a multi-agent framework to model the dynamic aspects of supply chains. This study created a library of software agents that capture generic supply chain processes and concepts, providing a modular and reusable framework for developing a wide range of realistic supply chain models. In contrast to Fox et al. [1993, 2001], which decomposed agents based on supply chain functions, Swaminathan et al. [1998] defined agents by considering the structural elements in the supply chain. The proposed framework in Swaminathan et al. [1998] allows models to address issues at both tactical (coordination) and strategic (configuration) levels. These early pioneering works laid the foundation for the subsequent development of multi-agent frameworks to address SCM challenges.
Since then, numerous applications of agent technology in SCM have emerged, which can be broadly categorised into four groups. The first category involves agent frameworks designed for generic or specific supply chains. These efforts aim to identify common characteristics of generic supply chains or individual industry-specific supply chain, providing a framework for constructing MAS for supply chain operations. For example, Sadeh et al. [2001] introduced a MASCOT architecture based on the blackboard architectural model [Erman et al., 1980]. This architecture providing an generic and reconfigurable agent framework for supply chain coordination and dynamic formation. Julka et al. [2002] defined a unified framework for modelling entire supply chain network in an object-oriented manner, with supply chain entities represented as agents and materials and information represented as objects. [Lou et al., 2004] proposed an agent architecture that consists of mediator and functional agents for agile SCM. Agent frameworks have also been applied in specific supply chains, such as fashion [Lo et al., 2008], wine [Kumar et al., 2013], outsourcing SME manufacturing [Kumar et al., 2013], and construction [Xue et al., 2005].
The second category focuses on utilising agent-based approaches to address dynamic supply chain formation and configuration. For example, Walsh and Wellman [1999] discussed the issues and models relevant to automated supply chain formation using an agent-based approach. Chen et al. [1999] introduced a MAS that utilises a set of performatives for pair-wise negotiation, enabling dynamic supply chain generation in an open environment through automated or semi-automated negotiation. Similar to Fox et al. [1993], this MAS consists of two types of agents: functional agents and information agents. Kim and Cho [2010] investigated the dynamic supply chain formation problem based on the Single Machine Earliness/Tardiness (SET) model and employed agent negotiation as a means to address this problem in a three-tier supply chain setting. Akanle and Zhang [2008] developed an agent-based model to manage customer demand over a period of time by optimising supply chain configuration. Meanwhile, Ameri and McArthur [2013] presented an agent-based framework, the Digital Manufacturing Market (DMM), for handling supply chain configuration. The DMM combined semantic web technology and agent technology to enable automated service discovery and matching, thereby facilitating dynamic configuration of a two-tier supply chain that only involves supplier and customers.
The third category relates to coordination and negotiation in SCM. While supply chain formation and configuration relate to the structural aspects of a supply chain, coordination and negotiation address its dynamical aspects [Swaminathan and Tayur, 2003]. Effective coordination among supply chain functions is essential for efficient operations. Barbuceanu and Fox [1996] tackled the coordination problem by explicitly representing interaction knowledge among agents and introduced COOL, a structured conversion-based communication language for supply chain coordination. Sadeh et al. [2001] developed the MASCOT architecture, which offers protocols for both horizontal and vertical coordination in supply chains. Xue et al. [2005] proposed an agent framework for construction supply chain coordination, based on multi-issue negotiation using multi-attribute utility theory (MAUT). Wong and Fang [2010] presented a contract-net-like multilateral and multi-issue protocol designed for handling buyers-seller bargaining and interaction negotiation in SCM. This protocol also adopts the MAUT approach to model utility functions during negotiations. Various general negotiation approaches suitable for SCM have been proposed, including the approach by Chen and Weiss [2015].
While the previous three categories focus on constructing agent-based SCM systems, the last category, which is main focus of most literature, aims to enhance specific functions or processes through MAS approaches, such as demand forecast [Liang and Huang, 2006], cost management [Fu and Fu, 2015], supplier selection [Ghadimi et al., 2018, Yu and Wong, 2015], scheduling [Aminzadegan et al., 2019], inventory planning and control [Ito and Mousavi Jahan Abadi, 2002, Kazemi et al., 2009, Kim et al., 2010], and risk and event management [Ahn et al., 2003, Bodendorf and Zimmermann, 2005, Giannakis and Louis, 2011, Bearzotti et al., 2012].
While there is a substantial body of literature on agent-based SCM (for a comprehensive review, see Xu et al. [2021] ), none of them focus on achieving the automation across the entire supply chain, including in both decision making and execution aspects. We in this paper aim to address this gap, developing methodology to design ASC systems using MAS approaches and presenting a technical framework for implementing a concrete A2SC prototype system.
3 Agent-based Autonomous Supply Chain (A2SC)
The adoption of technologies in the Fourth Industrial Revolution (Industry 4.0) [World Economic Forum, 2017] for transforming supply chains has gained significant attention [BlueYonder, 2020, Calatayud et al., 2019], with existing academic and industrial efforts mainly focusing on the application of digital technologies like AI and advanced robotics to specific functions, such as demand forecasting and pallet picking, rather than fully automating entire processes.
Automated functions are building blocks of fully automated processes, which, in turn, collectively form an entirely automated system. Within a supply chain, numerous functions require automation. Automating these specific functions needs multidisciplinary efforts and technological advancements, which is beyond of the scope of this paper. Instead, this paper focuses on integrating these “already-automated” functions and facilitating entire processes. Effective coordination mechanisms, both within individual business entities and across them, are crucial for achieving a seamless ASC. In this section, we describe MAS approaches, with a particular focus on the methodology for analysing and designing A2SC systems.
Agents are goal-directed computation entities that act autonomously based on their designated roles, while also communicating and coordinating with other agents as needed [Wooldridge and Jennings, 1995, Fox et al., 2001]. A Multi-Agent System (MAS) consist of multiple such agents, interacting to address complex problems that exceed the capacities or knowledge of any individual agent. MAS approaches are, therefore, well-suited for addressing problems involving: 1) Multiple (semi)-autonomous entities; 2) Distributed knowledge among many entities; 3) The need for entities to coordinate to accomplish specific tasks. Supply chains, which consist of a set of distributed business entities, exemplify such problems. Consequently, the MAS approach is an ideal choice for constructing ASC systems by coordinating automated functions and processes. The resulting ASCs are referred to as agent-based autonomous supply chains (A2SC).
3.1 Multi-Agent System Approach
Figure 1: Methodology for the analysis and design of A2SC systems.
Various methodologies have been proposed to analyse and design multi-agent systems, and the Gaia methodology [Wooldridge et al., 2000, Zambonelli et al., 2003] has gained recognition for its simplicity and effectiveness. However, Gaia has certain limitations; it lacks completeness in abstraction and thus has no capacity to identify organisational rules and model environment. Consequently, Gaia could only be used for closed communities with cooperating agents. To address these limitations, Juan et al. [2002] introduced ROADMAP, an extension of Gaia that incorporates additional models into its framework. Additionally, Zambonelli et al. [2003] extended Gaia by using organisational abstractions, resulting in an extended Gaia capable of analysing and designing complex open systems with self-interested agents. In contrast to ROADMAP, which is considered preliminary, this extended Gaia methodology is comprehensive and provides coherent integration with the original Gaia approach. We therefore chose to employ Gaia-based methodologies for the analysis and design of ASC systems, customising and adapting them to the specific context of supply chains, which are open systems and often involve self-interested agents.
Figure 1 illustrates the adapted methodology for the analysis and design of ASC systems, which is derived from the Gaia and ROADMAP methodologies. This methodology comprises five key stages, each of which involves specific primary tasks and output models, as outlined below:
  1. (1)
    Requirement Analysis: This stage determine the expectations of stakeholders for a new system under development or modification. The primary output artefacts are requirement specifications.
  2. (2)
    System Analysis: In this stage, an understanding of the system and its structure is developed. It includes the identification of primary roles, the environment in which they operate, and their interactions with other roles within this system. These roles serve as abstract constructs to conceptualise and understand the system. The primary outputs of this stage include roles model, interaction models, and an environment model.
  3. (3)
    Architectural Design: This stage defines the architectural aspects of these roles, including their acquaintance relation, organisational rules and structure.
  4. (4)
    Detailed Design: Detailed definition of agent models and services models are developed in this stage based on the roles, their interactions, and their organisation identified in the previous stages.
  5. (5)
    Implementation Design: While the previous stages mainly focus on analysing and designing the system’s structures without depending on specific implementation details, this stage is dedicated to designing elements that facilitate the system’s dynamics, particularly the flow of information through messaging. It involves three critical aspects related to agent interaction: knowledge representation, interaction protocols, and communication languages. Although these aspects do have relevance to implementation, they are included in this model as an independent design stage because the coherence of a MAS relies heavily on effective messaging. Other concrete system implementation aspects are excluded from this stage.
The methodology outlined above describes the two main phases that come before the development phase of implementing an ASC system. Each of these phrases comprises several stages, with each stage encompassing a set of key artefacts, as highlighted in blue in Figure 1. These artefacts can be broadly categorised into three groups based on their purpose: structural, social, and semantic. Structural artefacts, represented in light blue in Figure 1, are those models related to the composition and organisational aspects of a MAS. Social artefacts, highlighted in light green, pertain to the interactions and relationships among agents. Semantic artefacts, indicated in light violet, concern the understanding of the interaction content. These artefacts collectively contribute to the analysis and design of an ASC system, providing conceptual models for subsequent development phrase. In the following paragraphs, we delve into these artefacts by reorganising them into the four concrete issues when applying a MAS approach [Smith and Davis, 1981, Genesereth, 1997, Fox et al., 2001].
Decomposition and Organisation
In a MAS, as described in [Wooldridge and Jennings, 1995], multiple agents collaborate to achieve collective behaviours. These individual agents typically lack sufficient information or capability to independently solve the entire problem. Therefore, effective collaboration among these agents is essential to ensure the system operates coherently. To facilitate this collaboration, it is crucial to thoroughly decompose the problem and identify specific roles that agents should play in addressing the problem. Agents can be assigned one or more specific roles, defining their expected behaviours (or services) within the system. Additionally, it is important to organise these agents into a specific organisational structure that defines the topology of the interaction patterns and the control regime for activities within this organisation. The structural artefacts, which includes roles, agents and their services, and organisation as illustrated in Figure 1, are the main models for representing the composition of a MAS. While this issue tackles the static structural aspects of a MAS, the remaining three issues focus on its dynamics, including interactions and elements that facilitate communication.
Acquaintance and Interaction
Collaboration among agents is achieved through interaction. The design of agent interaction involves two key activities: 1) identification of the acquaintance relationship between agents, which determines whether two agents have a connection, i.e., communication paths, and 2) elaboration on the interaction process between these connected agents. The primary output of these two activities are social artefacts: acquaintance (or connection) models and interaction models. These models help identify any potential communication bottlenecks, ensuring the agents are loosely coupled. Well-designed interaction mechanisms can lower communication overhead, resulting in reduced data transmission and enhanced responsiveness.
Interaction Protocol and Communication Language
While the first two issues define the structure of a MAS, this issue deals with dynamics, specifically the flow of messages along the communication channels between its constituent agents. This includes establishing protocols that regulate interaction process and defining communication languages for representing messages. Some representative works in these areas include the Contract-Net protocol [Davis and Smith, 1983], and the English auction protocol [FIPA, 2001], and KQML [Finin et al., 1994], and FIPA-ACL [FIPA, 2002, Poslad, 2007].
Vocabulary and Knowledge Representation
With the first three issues addressed, the flow of messages is facilitated; however, message comprehension remains unaddressed. To ensure that agents understand the content of messages, they must use a standardised vocabulary with teams that consistently refer to objects, functions, and relations relevant to a particular application. This consistent vocabulary, along with a commonly understood knowledge representation, constitute the fourth issue that requires attention when applying a MAS approach. A vocabulary may encompass multiple ontologies used to describe the shared body of knowledge within a specific domain.
These four issues are essential when adopting an MAS approach. The first two issues pertain to the structural aspects (composition and organisation) of an MAS, whereas the last two relate to its dynamics, specifically, interaction and communication. Commonly communication languages and interaction protocols allow agents to freely join or leave the MAS, enabling an open, dynamical computational environment. These four issues are a concrete refinement of the methodology illustrated in Figure 1. Despite the differences between supply chains and traditional MAS scenarios, these four issues provide valuable guidance for the analysis and design of a A2SC system.
3.2 Analysis and Design of A2SC
An ASC, as described in [Xu et al., 2022], is an self-operating supply chain comprised of a set of structural entities. Each of these entities may represent individual business entities in a supply chain and can manage multiple automated processes. These structural entities are crucial components in the formation of an ASC and are responsible for managing the three flows within the supply chain. Efficient operation of an ASC requires effective coordination among these structural entities. However, their coordination is challenging due to their inherently self-interested, distributed, heterogeneous, yet cooperative nature. Fortunately, MAS approaches are well-suited for building open systems that involve both cooperative and self-interested entities.
Given the diverse range of automated sub-systems involved in an ASC, often using technologies from various domains, delving into the specific designs of these systems is neither necessary nor feasible within the scope of this article. Instead, we treat these automated systems as readily available building blocks for the assembly of an ASC. We then primarily focus on the integration of these structural entities and their internal processes, enabling seamless and autonomous collaboration. In this section, we discusses the analysis and design of A2SC system using the MAS methodology presented in the previous section.
In distributed problem-solving scenarios, such as distributed sensing and air traffic control, agents often share identical owners and are explicitly designed to work cooperatively towards a given goal. In such cases, problem decomposition typically involves organising roles and agents based on functionality or geographical location. However, in an ASC scenario, agents may represent different business entities that collaborate for mutual benefits, but they may also compete with others and thus have self-interests to consider. Moreover, ASCs are considered open systems that are collectively designed by stakeholders, in which stakeholders are free to join or leave this system.
In the context of A2SC, problem decomposition must take into account organisational boundaries. According to Xu et al. [2022], structural entities act as bridge points for business entities, allowing them to connect with external peers while managing their internal entities within a supply chain. Decomposing the supply chain into a set of structural entities, each represent multiple roles and potentially represented by an agent, aligns with the natural structure of the supply chain.
Despite operating at a higher level of abstraction, this decomposition approach somewhat aligns with with existing literature. Previous studies, such as Swaminathan et al. [1998], Fox et al. [2001], Akanle and Zhang [2008], identified two main categories of agents: structural and functional, based on their organisational and functional roles. Structural agents represent the organisational entities involved in a supply chain, including stakeholders like suppliers, retailers, third-party logistics (3PL) providers, distribution centres, and manufacturing plants. On the other hand, functional agents are responsible for specific functionalities in a supply chain, such as supplier selection, order placement, and logistics arrangement. A structural entity represent an conceptual organisation, thus it representative agent may manage multiple these structural and functional agents.
These decomposed agents collectively manage the operations of a supply chain through interactions. Agent interaction defines the potential interdependencies among agents in an MAS, thus influencing the MAS’s topology. Therefore, interactions among agents must be carefully designed to create a loosely-coupled, scalable MAS which agents focus more on computation rather than communication. In A2SCs, the management of the three flows — product, information, and financial — involves agents interacting with others to facilitate the inbound and outbound of movement of these flows. Developing appropriate interactions is crucial to ensuring the efficient coordination of these flows through a network of structural entity agents. During interactions, agents may exhibit two types of behaviours: proactive and reactive. Proactive behaviours involve agents initiating interaction to pursue their objectives, such as connecting other agents to request price data. Conversely, reactive behaviours are response to external events, such as handling incoming requests when receiving goods. Furthermore, interactions should be governed by protocols, which are the commonly recognised procedures or system of rules. These protocols manage various interaction processes, each regulating agents’ dialogue with others in different scenarios, such as supplier selection, procurement price negotiation, and advanced shipment notification.
Agents interact through messages; and in an A2SC, the entire system functions through the exchange of these messages. For an A2SC to function efficiently, these messages must be fully comprehended both syntactically and semantically by all participating agents. Metaphorically, agents in an A2SC community must “speak” a common language to enable clear and unambiguous communication. This common language should exhibit consistency in three critical aspects:
  • Message Structure: the way in which a message is organised or structured to convey information.
  • Representation of Message Content: the method used to represent the actual data contained in messages.
  • Ontologies: the shared vocabularies that describe the knowledge as a set of concepts within specific domains of a supply chain and the relationship between them.
These three aspects collectively standardise the messages in terms of their structure, content representation, and semantics. To achieve universally understood messages, they should be explicitly specified using common languages. Information flow among agents in an A2SC, including quotations, purchase orders, delivery status, and invoices, relies on effective messaging. Consequently, careful considerations on interaction protocols, languages, and message representation is crucial for the successful implementation of an A2SC.
4 A Case Study: an Autonomous Meat Supply Chain
Figure 2: Illustration of a simplified meat supply chain. The middle parts of this supply chain are the focus of our prototype.
The previous section introduces the A2SC and outlines the methodology for its analysis and design. In this section, we present a detailed case study of A2SC implementation to illustrate the development of an ASC prototype using an MAS approach. We selected the meat supply chain — a particular perishable foods supply chain — as the case for study. Perishable foods, unlike other goods, are prone to spoilage due to inadequate storage conditions. To ensure the successful delivery of perishable foods, it is crucial to maintain proper storage conditions and minimise transit times. Monitoring the environmental conditions of trucks transporting perishable foods and swiftly addressing unforeseen events, such as traffic congestion, is of paramount importance.
In this prototype, we address the procurement of meat, specifically implementing an end-to-end automated meat procurement process. This procurement process involves several key stakeholders in the meat supply chain. However, it is important to not that our representation of meat supply chain is simplified and simulated (Figure 2). This simplified version excludes minor participants, such as secondary processors, which does not exactly mirror the complexity of real-world meat supply chain operations. Figure 2 provides an overview of the main structure of this meat supply chain. Our prototype primarily focuses on the central components highlighted in grey.
Furthermore, this prototype includes logistics and third-party logistics (3PL) since they are responsible for transportation, which are essential to PFSCs. For simplification purposes, we omitted farmers and end customers. The core procurement process involving suppliers, wholesalers, and retailers is sufficient for illustrating an initial A2SC implementation, and including additional entities would lead to unnecessary redundancy.
4.1 Scenario: the Cambridge Meat Company
Figure 3: Illustration of the CMC scenario.
The prototype is based upon the following scenario:
The Cambridge Meat Company (CMC), a hypothetical company used for illustration purposes, specialises in wholesale meat procurement and supplies local restaurants. The CMC has established long-term contracts with its meat suppliers and aims to automate its wholesale and procurement process by incorporating disruptive information and communications technologies. Specifically, the company seeks automation in four key areas: 1) Selection of bids from suppliers; 2) Monitoring of the logistics process; 3) Adaptation to unforeseen events; 4) Evaluation of the quality of both the logistics service and the supplied products.
From this scenario, we can further analyse the requirements for the prototype, aiming to develop an end-to-end automated process for meat procurement and wholesale. A crucial requirement for this prototype is to ensure the smooth movement of physical flow and its corresponding information flow. Given that modern financial transactions primarily rely on digital information, essentially constituting a specialised from of information flow, this prototype does not encompass financial flow considerations. Additionally, this prototype take into account inventory management, continuously updating it as goods are received and dispatched. The following sections first extract requirements and then detail the design, implementation, and showcase of the A2SC prototype.
4.2 Process Design
(a) Replenishment
(b) Wholesale
Figure 4: Flowcharts illustrating the two main processes: replenishment and wholesale. Uppercase letters denote supply chain stakeholders: ‘S’ for supplier, ‘W’ for wholesaler, ‘R’ for retailer, ‘L’ for logistics, and ‘3PL’ for third-party logistics provider.
This prototype includes two primary processes: replenishment (the CMC procures meat from its suppliers to replenish its inventory) and wholesale (the retailer buys meat from the CMC). Figure 3 presents a three-dimensional diagram of the refined CMC scenario, with the two processes represented by the flow of blue and red arrows. In addition to the three main stakeholders (supplier, wholesaler, and retailer), logistics and 3PL, which provide logistics service, are included. Logistics service and inventory management, crucial for the flow of goods (meat), are considered in both processes. Both processes include logistics services, with sellers (supplier and wholesaler) responsible for delivery arrangements. Figure 4 illustrates the flowcharts of the these two processes, highlighting major steps. Different supply chain stakeholders are denoted by capital letters in brackets and are represented using different colours. The replenishment process (see Figure 3(a)) provides customers (wholesalers) with multiple delivery options, allowing them their preferred delivery service. However, the wholesale process (see Figure 3(b)) has a preselected delivery option by the wholesaler for their customers (retailers), making it less complex. Both processes involves multiple decision-makings points (denoted by diamonds), such as proposal evaluation, acceptance or rejection of proposals, and delivery option selection. Inventory-related functions, such as automatic inventory updates and replenishment, are integrated into both processes. Additionally, both processes incorporate logistics monitoring, which tracks real-time locations and ambient conditions of goods throughout delivery.
4.3 Roles, Agents, and Services
Figure 5: Agents and their primary services.
From the scenario described in Section 4.1 and its subsequent refinement processes (see Figures 3 and 4), we identified five distinct types of roles: suppliers, wholesalers, retailers, logistics providers, and 3PL provider, and each with their corresponding primary services, which are a coherent blocks of functions that roles are expected to accomplish. Each of these roles can be represented by individual agents, serving as the structural entities within the scope of their respective organisations. These agents and their main services are shown in Figure 5. A detailed description of each agent are as follows:
  1. (1)
    Wholesaler Agent: This agent represent the wholesaler, which is the central roles, namely the CMC, in this scenario. Its main services include managing the processes of procuring meat from supplier agents and wholesaling meat to retailer agents, and corresponding inventory management during these two functions. It also interacts with logistics agents to arrange delivery service for its customers. This agent serves as a central hub connecting upstream and downstream supply chain agents, facilitating the entire flows.
  2. (2)
    Supplier Agent: This agent is responsible for managing the process of supplying meat to wholesaler agents on behalf of a supplier. Similar to wholesaler agents, the supplier agent provides its customers with delivery service, which is outsourced to logistics agents. Therefore, this agent needs to coordinate with wholesale agents and logistics agents to fulfil purchase orders.
  3. (3)
    Retailer Agent: Representing local retailers, such as restaurants or local stores, this agent purchases meat from wholesaler agents. It interacts with the wholesaler agent to handle procurement on behalf of the retailer.
  4. (4)
    Logistics Agent: This agent, representing a logistics company, provides logistics services to supplier agents and wholesaler agents in the scenario. Logistics companies primarily manage overall logistics service but outsource fulfilment service to 3PL providers. Logistics agents need to coordinate with suppliers, wholesalers, and 3PL providers for delivery arrangements.
  5. (5)
    3PL Agent: This agent, acting on behalf of a 3PL provider, is responsible for providing delivery services to logistics companies, fulfilling delivery orders assigned by a logistics agent. It monitors the entire transportation process of the meat, providing logistics companies with real-time data on the truck’s geolocation and the meat’s ambient conditions for further use, such as fulfilment service evaluation.
In addition to the main services discussed above (indicated by solid-lined ellipses in Figure 5), these agents also offer information query services to other agents and include functions for (un)registration. Information query service provide these agents with the capability to connect with other agents to request specific information. Moreover, they have the ability to register their services into the system, making them ready for discovery and use. Conversely, they can unregister existing services, removing them from the running system and making them unavailable for discovery by other agents.
These five types of agents are the main stakeholders — structural entities — in this meat supply chain, and their interactions shape the dynamics of this supply chain. Additionally, we introduced an functional agent, known as the admin agent, responsible for managing agent-level administration services, including providing yellow page services. The following sections provide a more detailed description of the organisation and interactions among these agents.
4.4 Connection Mechanism
(a) Contract net protocol
(b) HTTP protocol
Figure 6: Interaction diagrams of the contract net protocol and the HTTP protocol. Note: The left is the FIPA contract net interaction protocol, which is a slight adaption of the original contract net protocol [Smith and Davis, 1981].
Before distributed agents can be organised together, the connections between them must be first enabled. We, thus, first introduce connection mechanisms for facilitating agent organisation and interactions. Agents in an MAS offer services to others or request services from others. Therefore, this leads to the challenge of how a service requester can find suitable service providers capable of executing a service. This challenge is known as the connection problem [Davis and Smith, 1983, Ameri and McArthur, 2013].
In this prototype, we devised two mechanisms for agent connection. The first is the Open Economic Framework (OEF)-based connection, in which two agents are linked through a third-party service known as the OEF. In this type of connection, a service requester submits a service query to the OEF, which responds by providing a list of potential service providers that match the query criteria. The service requester can then choose one ore more service providers from this list to establish further connections. The OEF-based connection is an indirect or mediated connection, providing a means for two unknown agents to discover each other through the assistance of a mediator. The second mechanism is the HTTP-based connection, in which two agents connect directly through the HTTP protocol, as explained in Section 4.6. In this approach, a service requester sends an HTTP request to a service provider via its URL and receives an HTTP response from that service provider. This represents a direct connection, which is well-suited for scenarios in which both agents are already known each other and can contact each other directly without the need for an intermediary.
4.5 Agent Organisation
Figure 7: Agent organisation, with OEF and HTTP-based connections denoted by solid and dashed lines, respectively.
Connection mechanisms enable agents to establish connections with others. To achieve coherence, connections among agents should be organised as certain topology. As described in Section 4.3, this prototype includes five types of structural agents and an functional agent. In designing its agent organisation, we have chosen not to include intermediary agents, such as blackboard, yellow page, and broker, for two primary reasons: a) this prototype involves a small number of agents, and b) the OEF is capable of providing efficient discovery and matchmaking services for agents. The resulting agent organisation is generally a collective of peers, as illustrated in Figure 7. These peers operate with independent control regimes, each focusing on specific stages of the three flows. The two connection mechanisms discussed in the previous section are used to connect these peers, denoted by solid and dashed lines, respectively. The admin agent can establish connections with all other agents via the HTTP connection, thereby maintaining an portal that provides an overview of the status of all agents in the system. The supplier, wholesaler, and retailer agents are linked by the OEF connection, enabling negotiations for replenishment or procurement. The HTTP connection becomes active between these agents if trust is established between them.
4.6 Agent Interactions, Language and Protocols
While the agents and their organisation define the structure of the A2SC system, this section discusses its dynamics, involving the interactions among these agents and the language and protocols that facilitate these interactions. As shown in Figure 4, the wholesale agent interacts with supplier agents in the replenishment process and with retailers during the wholesale process. Both processes require logistics services, thus involving two additional agents: logistics and 3PL provider agents. These agents interact by adhering to agreed-upon interaction protocols.
Figure 8: Agent interaction in the replenishment process.
Figure 9: Agent interaction in the wholesale process.
In this implementation, we incorporate two types of interaction protocols for facilitating interactions: the contract net protocol and the HTTP protocol. The contract net protocol is specifically designed for scenarios where a service requester needs to select suitable service providers capable of performing a specific task. This protocol often include multiple rounds of negotiation and information exchange, as illustrated in Figure 5(a). It is well-suited for managing complex interaction scenarios that may involve unknown participants, such as procurement processes. On the other hand, the HTTP protocol is employed for simpler interaction scenarios that typically involve require only a single round of interaction. As shown in Figure 5(b), a service requester, referred to as an HTTP client, can send a request to a service provider, an HTTP server, to either get data or post a task for execution. This protocol is most suitable for interactions between known participants who have already established a level of trust. For example, a logistics agent may directly assign a delivery task to one of its longstanding 3PL partners for fulfilment. Additionally, the HTTP protocol is appropriate for quickly retrieving information such as product wholesale price, delivery rates, and current traffic conditions.
The interactions among agent in the two processes are depicted in Figures 8 and 9, In both figures, both processes involve four types of agents, and these agents interact with each other through messages. The messages in the two figures are composed of two essential parts: performative and message body, delimited by a colon. The two parts respectively represent the communicative act (behaviour) and the content of the message. Interaction protocols define a set of performatives that may be used in the interaction. For instance, the contract net protocol includes performatives such as cfp (call for proposal), propose, and reject-proposal, while the HTTP protocol includes two basic performatives: request (which further has two types of requests: get and post) and response, as shown in Figures 8 and 9. Additionally, a message should include other metadata for unambiguous communication, such as sender, receiver, protocol, and ontology. The agent communication language defines the message format, and in this study, we we use FIPA-ACL as the language for agent communication.
Both processes involves multiple rounds of communication, sending and receiving messages to coordinate procurement and logistics service. They also include procedures to support decision-making, such as assess_order() for evaluating incoming procurement orders, and self-calls for executing specific tasks, such as monitor_delivery() for tracking delivery process. In the replenishment process, which offers multiple delivery options, there are additional rounds of interaction with the logistics agent, and subsequently with the 3PL agent, to request information about available delivery options. This is depicted in the right bottom area in Figures 8. Although the two processes are operationally separate and independent, the wholesale process may trigger the replenishment process automatically. For instance, when the wholesaler’s inventory reaches a predetermined low stock level, the replenishment process is initiated. In this sense, the two processes are interconnected, illustrating an end-to-end supply chain that facilitates the automated movement of “goods” from suppliers through intermediaries to customers.
4.7 Implementation
The previous sections provide an overview of the A2SC prototype’s design. In this section, we will first introduce the toolkit for developing this prototype, and then delve into its implementations.
4.7.1 Toolkit for Implementation
We used Python as the primary language for developing the backend of this prototype, while other languages such as HTML and JavaScript were used for creating the frontend interfaces. To streamline the development process, we leveraged a toolkit consisting of the following tools and frameworks:
  • AEA Framework: The AEA (Autonomous Economic Agent) framework [Minarsch et al., 2021] from Fetch.ai  [^5] https://fetch.ai/ technology stack is a Python-based toolkit for constructing autonomous agents known as AEAs. These AEAs operate independently on behalf of their owners and autonomously execute actions to achieve predefined objectives. We used this framework to develop the agent network in the prototype.
  • OEF (Open Economic Framework): The OEF serves as a decentralised communication, search and discovery system for AEAs, providing protocols, languages, connection mechanisms that facilitate agents search, interaction discovery, and engagement within established social norms. It establishes a virtual environment where the inhabitants, AEAs, can operate and interact. In Fetch.ai technology stack, the AEA framework and the OEF represent the top two layers, positioned above the smart ledger layer, as depicted in Figures LABEL:fig:fetch_three_layers. For this implementation, we exclusively utilised the AEA framework and the OEF to develop the prototype, as they are capable of functioning independently without the smart ledger.
  • REST and OpenAPI: REST (REpresentational State Transfer) Fielding [2000] is an architectural style for providing interface standards between decoupled components on the web. RESTful APIs are typically based on HTTP methods to access resources via URL-encoded parameters and the use of JSON or XML for data transmission. All agents in the prototype were equipped with RESTful APIs, enabling them communicate with each other without the OEF and access external resources. However, these agents need to expose their APIs to other in a standard manner. The OpenAPI specification (OAS) [OpenAPI Initiative, 2021] was therefore adopted for describing these interfaces. The OAS defines a standard, programming language-agnostic interface description for RESTful APIs, allowing both humans and machines to discover and comprehend a web service without requiring access to source code and extra documentation. In this prototype, we define interface description files (.yaml) by adhering to the OAS.
  • WebSockets: WebSockets, which enable bi-directional, full-duplex communication, are well-suited for frequently updated or real-time applications. We thus used WebSockets [Fette and Melnikov, 2011] for services such as notification, monitoring, and agent chatting. The above tools are relevant to backend development and architectural design. We will describe the tools/frameworks used for frontend development in this following paragraphs.
  • Django and Django Channels: Django  [^6] https://www.djangoproject.com/ is a Python-based, fast, and easy-to-use web framework that follows the model-template-views (MTV) architectural pattern. It emphasises reusability and pluggability of components, which can serve as a hub to integrate heterogeneous components such as AEAs. Django, while versatile, is not well-suited for protocols require long-running connections, such as WebSockets, MQTT, and chatbots. Channels  [^7] https://channels.readthedocs.io/en/stable/ extends Django’s capabilities to handle both HTTP and real-time applications. We used Django and Channels for developing web-based interfaces and integrating various UI elements into a browser-accessible portal.
  • Leaflet and OpenStreetMap: Leaflet  [^8] https://leafletjs.com/ is an open-source JavaScript library designed for building interactive maps. It offers lightweight, user-friendly, extendable, and cross-platform map creation capabilities. OpenStreetMap (OSM)  [^9] https://www.openstreetmap.org, rather than other map services like Google Maps, was used as the map services due to its free of charge and sufficient effectiveness. We utilised Leaflet in conjunction with OSM to provide map service for monitoring logistics activities in this prototype.
  • Chart.js: Chart.js is a free, open-source, and simple-to-use JavaScript library for creating HTML-based charts. We used Chart.js  [^10] https://www.chartjs.org/ to visualise real-time IoT data related to the ambient conditions of goods in this prototype.
In addition to these major tools, we incorporated complementary tools and frameworks such as Bootstrap  [^11] https://getbootstrap.com/, JQuery  [^12] https://jquery.com/, and Ajax. This toolkit instrumental in enhancing communication, interaction, and visualisation functionalities for both backend agent network and frontend interfaces during the development of the A2SC prototype.
4.7.2 System Architecture
Figure 10: The system architecture of the prototype, with WS stands for the WebSocket protocol.
The A2SC prototype primarily consists of two main parts: an agent network and an interface web application. We implemented a microservice architecture to integrate these components using RESTful APIs and WebSockets, as depicted in Figure 10. The agent network is composed of a collection of AEAs, with each representing distinct organisations or functions. These AEAs operate within an “environment”, specifically the OEF, and collaborate cohesively to pursue their respective objectives. Additionally, they may also access storage services to persist data and utilise sensors to sense the environment. The interface application, powered by Django and Channels, provides a web-based portal to the prototype. The interface comprises a client with a collection of UI components, which we will further described in the next section. Additionally, it includes a server that handles HTTP and WebSockets requests through its Views and Consumers modules, respectively. Specifically, it handles user input and provides real-time visualisation of the agent network’s state through RESTful APIs and/or the WebSocket protocol. This architectural design enables a loose coupling between the agent network, which may be distributed over many machines, and the interfaces, allowing them to maintain independence and extensibility.
4.7.3 Interface
Based on the predefined scenario and the system architecture, specifically the left part presented in Figure 10, we developed a single-page interface for this prototype. The web-based interface is constructed using a Channels-extended Django framework, enabling access through a web browser. The interface consists of four main panels, each serving a specific purpose:
  1. 1)
    Ordering: This panel provides users with an interface for placing orders to purchase products.
  2. 2)
    Logistics Monitoring: This panel visualise the real-time track and trace of the vehicle fulfilling the delivery.
  3. 3)
    Streaming Data: This panel displays the real-time charts of the ambient conditions of the products during transportation, including temperature, humidity, and lightness.
  4. 4)
    Agent Chat Room: This panel shows the negotiation process among the agents, including message exchanges between them.
These panels interact with the agent network using either HTTP or WebSockets. We use Ajax for the ordering panel and WebSockets for the other other panels to dynamically update the current interface with new data from the backend systems, enhancing use experience by eliminating the need to load entire new pages.
4.7.4 Data Preparation
As previously described, this prototype involves real-time data collection to monitor meat delivery processes. Instruments such as sensors and GPS trackers are required to capture the ambient condition of the meat and the location of its transportation vehicle. This data needs to be continually collected during the delivery process. However, due to limited experimental condition, it is not feasible to actually drive a vehicle with necessary data collection instruments and refrigerated containers to transport meat from one place to another while establishing a wireless communication link with remote agents. Moreover, using a realistic setting for this prototype would be uneconomical and unnecessary, particularly in light of the primary objective of this study, which is to showcase the technical methodology rather than achieving a high level of technical readiness.
To address this, we used pre-collected IoT and GPS data to simulate real-time data collection during the delivery process. These data were collected through the following steps:
  1. 1)
    Select multiple locations in Cambridge, UK, as the locations for the suppliers, the wholesaler, and the retailer.
  2. 2)
    Install a UbiBot all-in-one sensor  [^13] https://www.ubibot.com/ubibot-ws1/ and a smartphone with a GPS recording App (myTracks [^14] https://apps.apple.com/us/app/mytracks-the-gps-logger/id358697908) on a bike.
  3. 3)
    Ride the bike from one selected location to another, capturing the data about the ambient condition of the bike and its location once every five seconds over this journey.
  4. 4)
    Calibrate the collected data and save them as CSV files for subsequent use in prototype development.
The collected data consists of a series of IoT (temperature, humidity, and light) and geolocation data (longitude, latitude, and elevation) points indexed in chronological order, essentially forming time series data. We utilised these CSV files to simulate real-time data capture during the delivery monitoring.
4.7.5 Implementation Details
Following the design described previously, we implemented this prototype using Python and the toolkit presented in Section 4.7.1. In this section, we will delve into the specifics of the implementation.
We adopted a loosely coupled architecture in which the Django-powered interface and the AEA-powered agent network were developed independently. These two components were integrated into a unified system through the use of RESTful APIs and the WebSocket protocol. In total, we implemented seven distinct agents within the agent network, comprising one supplier, one wholesaler, two retailers, one logistics, two 3PL providers, and one admin. These agents were individually developed using the AEA framework and collaborate by exchanging messages following predefined interaction protocols.
The AEA framework provides a base structure for agents, resulting in the directory structure employed for agent development in this prototype. The agent directory structure is illustrated in Figure 11. As shown, an agent is composed of four main packages for further development: connections, contracts, protocols, and skills. These packages are responsible for implementing agent-specific business logic. In out implementation, we used built-in implementations for connections, and protocols. Since this prototype does not include any features related to blockchain, we did not consider contracts, which typically contains wrappers for smart contract logic. Our primary focus was on skills, which define the specific services and capabilities of an agent. An agent may possess multiple skills, each tailored to handle interactions with other agents under specific scenarios. Skills are divided into types of actions: reactive and proactive, implemented through a set of handlers and behaviours, respectively. Handlers are responsible for processing incoming messages following a specified protocol, constituting an agent’s reactive behaviour. Behaviours encompass self-initiated actions driven by internal logic rather than external messages. Additionally, a skill incorporates data models for maintaining shared state within a skill, including description of the services that an agent with this skill can offer.
To manage CPU-bound or long-running executions, the framework provides the capability to create tasks within handlers and behaviours. We utilised this capability to handle long-executing tasks such as monitoring the delivery process. Decision-making logic, such as assessing a proposal, is organised into a strategy file within a skill. Additionally, a skill may contain files related to agent dialogue management (dialogues), database connections (db_communications), and miscellaneous utilities (helpers). An agent’s capabilities are configurable through YAML files, allowing tasks such as configuring a skill or specifying a RESTful service. For a more comprehensive understanding of the AEA framework, refer to the work by Minarsch et al. [2021].
Figure 11: A visual representation depicting the directory structure of agents within this prototype.
The implemented agents have access two types of data sources: databases and IoT sensors. Agents such as supplier, wholesaler, and retailer manage their own inventory systems, which were simulated by updating their respective inventory table when products are inbound or outbound. PostgreSQL was used as the database system to create these tables, which contains a list of products and their essential details, such as product name and quantity. The 3PL agents required access to IoT sensors to for delivery monitoring. This monitoring process was simulated by periodically read the data that was pre-captured using the method described in Section 4.7.4.
The web interface for this prototype was built using Django and its plugin Channels, as the underlying technologies. WebSockets and RESTful APIs were used to exchange data and control between the web interface and the agent network, effectively connecting the frontend and backend of the system. Real-time data, including vehicle location, agent messages, and system notifications, were transmitted to the interface using WebSockets, in which consumers was created to handle these data. These data were then rendered using appropriate frontend techniques. For instance, Chart.js was used for visualising IoT data, while OSM and leaflet were employed for tracking and monitoring the delivery vehicle. The agent chat room was developed from scratch without relying on specific libraries. The ordering panel was implemented using Ajax, a set of client-side techniques that enable web applications to send and receive from a server (i.e., an agent) asynchronously without interfering with the display and behaviour of the existing page. To enhance the layout and styling of the interface, Bootstrap and JQuery were utilised. Further details about the tools used in developing the prototype are provided in Section 4.7.1.
4.8 Showcase
Figure 12: A screenshot of the startup interface of the prototype. This interface consists of four areas: the top left panel (ordering), the top middle panel (map for delivery monitoring), the right panel (agent chat room), and the bottom panel (IoT data visualisation). All of areas are is in their initial states, with empty content or default values.
Once the OEF, all agents, the Django built-in web server, and the PostgreSQL service are started, the prototype is ready for use. Upon initialisation, each agent will execute a registration behaviour. This behaviour is a one-shot proactive action that involves sending a message to the OEF for registering their service. This registration allows them to be discovered by other agents in the system.
The prototype can be accessed via a web browser. Figure 12 shows the startup interface of the prototype, which includes the four panels as described in Section 4.7.3. Initially, no data is displayed on the interface besides the default values in the ordering panel. These default values include the default quantity and unit price of the meat to purchase, default performance and location requirements for vendors, and the default scenario. In this section, we will showcase this A2SC prototype, focusing on the two processes: procurement and replenishment.
At the beginning of the execution,the wholesaler’s inventory is empty. To prevent running out of stock when the retailer procures meat from the wholesaler, we initiate a replenishment process. We select the ‘replenish’ scenario and click the ‘Launch’ button with other default values. Clicking the ‘Launch’ button triggers the replenishment behaviour of the wholesaler agent. The wholesaler agent first establish connection with the OEF and sends it a query to search suitable suppliers. In response, the OEF returns a message containing a list of supplier agents that meat the query conditions back to the sender. This process is essentially a matchmaking step, matching a service requester with a suitable provider. The wholesaler can then start dialogues with the discovered suppliers, negotiating with a selected supplier for the purchase of products. The logistic agent also participates in this dialogue, and the supplier requests available delivery options from this agent. When all parties agree on the deal, the logistics agent randomly selects one 3PL from a set of eligible 3PL agents to fulfil the delivery. The delivery process begins, and the 3PL agent is responsible for monitoring and managing the delivery.
The resulting interfaces during the replenishment process are shown in Figures 12(a) and 12(b). Figure 12(a) shows the interface at the moment when dialogue is completed, and the delivery is just starting. At this stage, no IoT data has been collected yet, so there is no data displayed on the charts below the map. The notifications at the top right corner of the interface, as presented in Figure 12(a), provide key progress updates, such as indicating that the dialogue is done and the delivery has commenced. Figure 12(b) presents a screenshot of the interface during the delivery process. The delivery is visually represented by a moving vehicle with a blue-line trace on the map, along with an overlaid tracking number and a blinking red circle. The ambient condition inside the delivery vehicle are also captured and visualised on the charts.
When the product has been successfully delivered, the relevant data collected during the delivery journey are gathered, and a summary report is generated using a data exploration tool called Pandas profiling  [^15] https://pandas-profiling.ydata.ai/. This summary report can be accessed by clicking the ‘Report’ dropdown button or the stacked icons at the top right corner of the map. The aim of this summary report is to demonstrate the potential use of the data for evaluating the delivery service, and therefore, a quick exploratory data analysis tool was adopted for report generation. The corresponding interface screenshot is shown in Figure 12(c). A retailer’s automated procurement is similar to the wholesaler’s replenishment, as described in Section 4.2. When the procurement is complete, the entire trace of the delivery vehicle is shown on the map, as depicted in Figure 12(d).
The dialogues within this process occur through messages, which are displayed in the agent chat room area. Excerpts of agent chat messages are presented in Figures 14 As shown in Figures 14, messages have four properties: sender, recipient, performative, and message body (content). The message body contains data associated with the message, such as procurement orders, receipts, and delivery notifications. These data are represented using JSON (or Javascript Object Notation). JSON is a human-readable data interchange format suitable for storing and transmitting data in web applications with servers. A screen recording of this prototype, providing more detailed demonstrations, is accessible via this link  [^16] https://www.youtube.com/watch?v=WFcLAif-cHo.
5 Discussion and Limitations
The prototype provides a concrete system to showcase the implementation of an ASC using a MAS approach. However, it is worth noting that this prototype may not strictly qualify as a complete ASC. Instead, it is an experimental proof of concept aligned with the MISSI model — a conceptual model described in Xu et al. [2022]. In this implementation, the main is placed on the two central layers of the model: interconnection and integration. The MAS approach was adopted as the overarching architecture to facilitate the interconnection and integration of decentralised business entities along the supply chain, creating a cohesive whole. The coherence of this whole is achieved through interaction enabled by message flows among the participating agents. To facilitate effective interactions, we need to address the following three key aspects:
  • Representation: This involves formally describing the current status or knowledge of an issue, request, or response. It is about how we represent data, such as bids or events, in a structured and machine-readable manner.
  • Decision-making: This aspect concerns the evaluation of incoming requests and the process of making appropriate responses. It is about defining rules and criteria for determining the best course of action in response to various issues.
  • Social: The focus is on establishing shared standards of acceptable behaviour, protocols, and language that enable smooth interactions with others. It involves defining the norms and conventions that govern how agents communicate and collaborate within the system.
In this prototype, we have indeed considered these three aspects, but our implementation is relatively basic and limited in scope. For representation, we primarily addressed how to represent the service that agents can offer and some data objects in the scenario, such as purchase orders and receipts. Regarding decision-making, our approach was rule-based and relatively simplistic. For example, the strategy of a wholesaler for accepting or rejecting a purchase order relies on a simple rule: it accepts the order if it has sufficient product to supply and the purchase price is higher than expected; otherwise, it rejects the order. We utilised the contract-net protocol to govern interactions and coordinate task allocation. However, it is important to note that this protocol alone may not suffice for handling complex real-world interactions that involve intricate negotiations. In terms of social interactions, our prototype does not incorporate formal languages for knowledge representation and social protocols. This prototype represents an initial step toward building a coherent A2SC system, which mainly focuses on automating a series of functions through data exchange. We underscore the need for carefully designed formal or even natural languages for knowledge representation and social protocols.
According to the autonomy manifold presented in Xu et al. [2022], this prototype falls within the automation-skewed region. However, both the intelligence and automation levels of the prototype are relatively low. The intelligence dimension primarily concerns decision-making capabilities. In this prototype, decision making is rudimentary, relying on a set of simple rules that consider only a few factors. To address real-world challenges in the supply chain domain, the agents need to be equipped with more sophisticated decision-making capabilities, enabling them to make informed decisions and propose appropriate courses of actions. The automation dimension of this prototype mainly focuses on automating data flow and process execution. This automation is facilitated through interaction among AEAs, which can request services or respond to requests, ultimately streamlining the overall process. However, this automation is limited to known data transmission paths, and the prototype neglects important aspects such as financial and product flows, which are integral to a functional supply chain. For an effective ASC system, it is crucial to address and enhance both the intelligence and automation dimensions to handle a broader range of scenarios and challenges.
The prototype includes two distinct processes: replenishment and wholesale. Both processes can be initiated manually by a user who clicks the ‘Launch’ button after completing an order form with the desired values. While the objective is to achieve level 2 autonomy, which implies process automation, the prototype falls short by only partially automating its functions. Automation is limited to ordering and order fulfilment in a controlled simulation environment, with financial and physical flows remaining unrealised. Additionally, SCM functions such as planning are absent from this prototype, which are essential for an ASC reach higher autonomy levels.
This prototype involves approximately ten stakeholders, which is significantly fewer than what is typically encountered in real-world supply chain scenarios. Managing a large number of agents can introduce complexities in various aspects, including architectural scalability, protocol design, and managerial challenges. In reality, ASCs must address these complexities to create a dynamic and adaptable society of living agents capable of self-adjusting to accommodate changes and uncertainties.
Furthermore, as highlighted in Xu et al. [2022], the development of ASCs involves the consideration of critical aspects that extend beyond technological efforts. These aspects include cyber security, data security and privacy, trustworthiness, and platform neutrality. Effectively addressing these dimensions need collaborative efforts from all participants within the ASC ecosystem. As an initial proof-of-concept, this implementation may not have fully addressed all of these aspects. However, its adoption of an MAS approach for integrating distributed entites represents a promising step toward addressing platform neutrality in the broader ASC domain.
6 Conclusion and Future Work
This paper focuses on the implementation of an autonomous supply chain (ASC) system, which is an emerging concept within the field of SCM. Specifically, we introduce a model for ASC implementation known as the agent-based autonomous supply chain (A2SC). A2SC is an ASC system that is realised through the utilisation of a multi-agent system (MAS) approach.
We have presented a methodology designed for the analysis and design of A2SC systems, which is adapted from a classic agent-oriented analysis and design framework, the Gaia methodology [Wooldridge et al., 2000]. We have provided a detailed concrete case study, the autonomous meat supply chain, demonstrating the practical implementation of A2SC systems using our proposed methodology. Additionally, we presented a system architecture and a toolkit for developing A2SCs. While the architecture and tooklit are specifically designed for this case, they can offer valuable guidelines for realising other similar decentralised systems.
As discussed in the previous section, this work represents an initial attempt to implement an ASC system, and it comes with certain limitations. Our future endeavours will primarily focus on addressing these limitations, which include 1) development of appropriate languages for agent communication and content representation within A2SCs, 2) designing or introducing protocols capable of handling more complex scenarios, such as bargaining and multi-issue negotiation, 3) scaling up the problem, including expanding the number of stakeholders, and 4) conducting experiments in less restricted setting. Furthermore, we plan to explore the potential applications of emerging technologies and concepts within ASCs, such as Digital Twin (DT) and artificial general intelligence (AGI). A2SCs can be viewed as a unique form of supply chain digital twins, serving as not only virtual representations but also control centres for actual supply chains. This perspective positions DT technologies and related concepts as promising avenues for advancing ASC development. Additionally, the recent tremendous progress witnessed in AGI systems, particularly in the domain of Large Language Models such as ChatGPT [OpenAI, 2023] and Llama [Touvron et al., 2023], offer intriguing opportunities for A2SCs. These advancements enable the creation of intelligent agents capable of making human-like decisions — a pivotal capability for effectively managing various supply chain tasks in A2SCs. We will thus explore the potential of AGI technologies in development of ASC systems.
In conclusion, this study, to the best of our knowledge, is the first work to report on a technical implementation case of an ASC system. It demonstrates a promising pathway for the development of an effective ASC system. The methodology, system architecture, toolkit for implementation, and agent roles and services model presented in this paper shed light on the future research and technical implementation of ASC systems in various scenarios.
Acknowledgement
This work was funded by the Research England’s Connecting Capability Fund (grant number: CCf18-7157): Promoting the Internet of Things via Collaboration between HEIs and Industry (Pitch-In) and the EPSRC Connected Everything Network Plus under grant EP/S036113/1.
Appendix A Additional Figures
(a) Replenishment: dialogue completed, delivery initiated. Relevant system notifications appear in the top right corner of the interface. On the map, you can see the locations of the supplier and wholesaler, which are the source and destination for the delivery with the tracking number: Hermes1594666109633. The ongoing dialogues between the agents are displayed in the agent chat room on the right.
(b) Replenishment: delivery in process. The vehicle is current in transit, transporting the meat to its destination. Its real-time location is tracked and traced, represented by a blinking circle with a blue solid line on the map. Additionally, the real-time ambient conditions of the meat, including temperature, humidity, and lightness, are visualised through the charts on the bottom area.
(c) Replenishment: delivery completed, summary report available. Once the products have been successfully delivered, a summary report generated from the collected the IoT data during the delivery journey is available to access. To view the report, simply click by the ‘Report’ button located in the top right corner of the interface. A portion of the summary report will appear in a scrollable pop-up window at the centre of the interface.
(d) Replenishment and wholesale (procurement): delivery of both processes completed. When the wholesale process is complete, the entire delivery route is displayed on the map, and a complete list of agent dialogue messages are shown in the agent chat room. Their respective delivery summary reports can be accessed by either clicking the Report button or the light grey icon on the top right corner of the map.
Figure 13: Example screenshots of the prototype during the two processes: replenishment and wholesale (procurement).
Figure 14: Agent chat message excerpts: (a), (b) and (c) are messages during the replenishment process; and (d) is for wholesale.